Let X be a discrete random variable that follows a binomial distribution with n = 11 and probability of success p = 0.31.

What is P(X=2)?

Round your response to at least 3 decimal places.
Answer #1

Solution

Given that ,

p = 0.31

1 - p = 1 - 0.31 = 0.69

n = 11

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 2 ) = (11C 2) * (0.31)2 * (0.69)9

= 0.18738

Probability = 0.187